# Probing the accretion geometry of black hole X-ray binaries: A multi-mission spectro-polarimetric and timing study

### v1

@@ 4.1.2 Power Density Spectra
We investigate the power density spectra (PDS) of all sources in different spectral states using NICER observations (see Table 1) in $0.5-10$ keV energy range. In addition, PDS properties in hard X-rays ($>10$ keV) are also examined using available NuSTAR and AstroSat observations of the sources. We generate 0.01 s time binned light curves and compute the PDS up to Nyquist frequency of $50$ Hz with $1024$ newbins per interval using powspec inside HEASOFT V6.34. The individual power spectra are averaged out to get the resultant PDS in units of $\rm rms^{2}/Hz$ which is geometrically rebinned in the frequency space with a factor of $1.03$ (see Belloni et al., 2002, 2005; Sreehari et al., 2019b; Majumder et al., 2022). In the upper panels of Fig. 2, we present the PDS of the respective sources in different spectral states. We model each PDS with a combination of constant and multiple Lorentzian components. Further, we compute the total percentage rms amplitude ($rms_{\rm tot}\%$) as a measure of variability in $0.1-50$ Hz and tabulate it in Table 3.
@@
Interestingly, we observe that the PDS in several spectral states exhibit distinct variable features. In particular, PDS of Cyg X11-1- 1 in the LHS show significant variability with $rms_{\rm tot}\%\sim 31.1-52.4$ ($0.5-10$ keV) and a clear power spectral break at $\sim 1.5$ Hz with NICER observations. Further, power spectral properties in hard X-rays ($3-80$ keV) with quasi-simultaneous AstroSat observations exhibit similar break frequencies ($\sim 1.5$ Hz) with $rms_{\rm tot}\%\sim 37$ in LHS (see Table 1 and Fig. 2a). We also observe marginal (significant) PDS variability of $rms_{\rm tot}\%\sim 9.1$ ($33$) with low (high) energy NICER (AstroSat) observations of Cyg X11-1- 1 in the HSS. Note that similar variability properties including power spectral breaks were also reported by Kushwaha et al. (2021) in LHS and HSS of Cyg X11-1- 1.
@@
For Swift J1727.816131727.816131727.8-16131727.8 - 1613, the presence of type-C QPO feature at $\sim 1.4$ Hz is observed during HIMS (epoch 1, hereafter E1) in $0.5-100$ keV with NICER, NuSTAR and AstroSat quasi-simultaneous observations (see also Nandi et al. 2024; Liao et al. 2024). The evolution of the QPO frequency up to $\sim 6.7$ Hz is seen with NICER in the later epochs (E2-E5). However, a power distribution, similar to Cyg X11-1- 1 (LHS), is observed in the LHS of Swift J1727.816131727.816131727.8-16131727.8 - 1613 without the presence of power spectral break, whereas marginal variability with $rms_{\rm tot}\%\sim 1.4$ is noticed during HSS (E6-E7).
@@
The remaining sources (4U $1630-47$, LMC X11-1- 1, LMC X33-3- 3, 4U $1957+115$, Cyg X33-3- 3 and Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147) in SIMS and HSS display weak variability signatures mostly represented by constant noise distribution and occasionally marginal power variation at lower frequencies. Conversely, during the recurrent outburst of GX $339-4$, a sharp type-B QPO peak at $\sim 4.5$ Hz ($rms_{\rm QPO}\%\sim 10.2$) and harmonic feature at $\sim 9.5$ Hz along with weak broadband noise ($rms_{\rm tot}\%\sim 15.4$) are observed in SIMS (see Fig. 2 and Aneesha et al. 2024). Note that, the QPO features with enhanced $rms_{\rm QPO}\%$ from $3.76$ ($0.5-10$ keV) to $12.6$ ($3-80$ keV) are observed with NICER and AstroSat, respectively. For IGR J17091362417091362417091-362417091 - 3624, a sharp type-C QPO at $\sim 0.2$ Hz ($rms_{\rm QPO}\%\sim 15$) along with a weak harmonic at $\sim 0.4$ Hz are observed during the onset phase of the recent outburst (see Fig. 2).
@@ 4.2.1 Spectral Modeling
Figure 3: Variation of the data to model ratio of the HSS/HIMS spectra of Cyg X11-1- 1, GX $339-4$ and Swift J1727.816131727.816131727.8-16131727.8 - 1613 fitted with model combination M1 and M3. The improvement in the respective fits by incorporating the reflection model component over M1 is depicted in panels (b), (d) and (f). See the text for details.
@@ 4.2.2 Alternative Spectral Models
We aim to explore the possible alternative disc-corona geometries in explaining the high-energy tail of the observed spectra. To achieve this, we use the comptt (Titarchuk, 1994) component as a proxy for thcomp in modeling thermal Comptonization. Note that, the ‘geometry switch’ selected inside comptt model as ‘$\leq 1$’ enables a disc-like geometry for the Comptonizing corona. Considering this, we proceed in the spectral fitting of all sources with the model combination constant×$\times$×Tbabs×$\times$(diskbb + comptt) (hereafter M2). Interestingly, we find that the model M2 provides a fit equally as good as M1 for all observations under consideration except a few observations of Cyg X11-1- 1 (E6 and E7), Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E4 and E5) and GX $339-4$ (E1). This leads to a possible degeneracy between the selected models in describing the observed spectra and the geometry of the Comptonizing region.
@@ 4.2.3 Spectral Properties
Figure 4: Confidence contours ($3\sigma$) in the PD–PA plane obtained from the model-independent polarization measurements using $2-8$ keV IXPE data, combining all DUs across various spectral states of the sources. The epoch of each observation is indicated in parentheses next to the corresponding spectral state. For Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, LMC X11-1- 1 and MAXI J174429417442941744-2941744 - 294, null-detection of polarization is observed, whereas all other sources exhibit significant polarization. The inset highlights a zoomed-in region of the PD–PA space for better clarity. See the text for details.
@@
We find that the effects of Comptonization are reasonably high during the LHS of Cyg X11-1- 1 with covering fraction $cov_{\rm frac}\sim 0.61_{-0.04}^{+0.04}-0.77_{-0.02}^{+0.02}$, where $cov_{\rm frac}$ measures the fraction of seed photons that are being up-scattered by thermal electrons of the corona. In addition, a high Comptonized flux contribution of $73-80\%$ is also observed in the spectra (see Table 3). Indeed, such a high $cov_{\rm frac}$ results in low photon index ($\Gamma_{\rm th}\lesssim 1.64_{-0.02}^{+0.02}$) as observed during the LHS observations of Cyg X11-1- 1. In the HIMS and LHS of Swift J1727.816131727.816131727.8-16131727.8 - 1613, we find $cov_{\rm frac}\sim 0.35_{-0.01}^{+0.01}-0.82_{-0.02}^{+0.02}$ and $\Gamma_{\rm th}\sim 1.71_{-0.02}^{+0.03}-2.11_{-0.04}^{+0.05}$ with up to $\sim 64\%$ Comptonized emission. Similarly, LHS of IGR J17091362417091362417091-362417091 - 3624 which is dominated by the Comptonized emission ($\sim 87\%$ flux contribution) yields harder ($\Gamma_{\rm th}\sim 1.62_{-0.01}^{+0.02}$) spectral characteristics. On the other hand, the HSS/SIMS spectra of Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613 results in low Comptonization with $cov_{\rm frac}$ in the range $0.005_{-0.002}^{+0.002}-0.38_{-0.03}^{+0.04}$ and thermal disc emission with $39-98\%$ flux contribution along with a steeper spectral index ($\Gamma_{\rm th}\sim 2.32_{-0.04}^{+0.06}-5.14_{-0.39}^{+0.36}$). Moreover, HSS with $76-98\%$ disc contribution are observed for LMC X11-1- 1, 4U $1630-47$, LMC X33-3- 3 and 4U $1957+115$ with negligible Comptonized emission ($cov_{\rm frac}<0.1$) as shown in Table 3. We find that GX $339-4$ (Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147) exhibits intermediate spectral characteristics with $cov_{\rm frac}\sim 0.34$ ($0.21$) and $\Gamma_{\rm th}\sim 2$ ($2.5$) with $\sim 50\%$ ($78\%$) disc emission during SIMS (HSS).
@@
The reflection spectroscopy of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613 and GX $339-4$ with model M3 during the HSS/SIMS results in disc ionization parameter ($\log\xi$) in the range of $3.31_{-0.05}^{+0.05}-4.29_{-0.21}^{+0.14}$ erg cm $\rm s^{-1}$, iron abundance ($A_{\rm fe}$) between $\sim 4.61_{-0.31}^{+0.36}$ and $6.35_{-1.46}^{+2.08}$ $A_{\odot}$, and reflection fraction ($R_{\rm f}$) ranging $\sim 0.59_{-0.11}^{+0.05}$ to $3.16_{-0.33}^{+0.35}$. Here, $A_{\odot}$ represents the solar iron abundance. The best-fitted parameters of the relxill component obtained from model M3 are tabulated in Table 4, whereas the parameters associated with other spectral components (diskbb and thcomp) of M3 are presented in Table 3 for the respective sources. Notably, a high ionization ($\log\xi\sim 4$) observed in HSS/SIMS generally suggests the reflection of significant coronal emission producing strong reflection signatures.
@@
In addition, we find that the model component relxilllp (Dauser et al., 2013) also delineates the reflection features assuming the primary source emissivity profile from a lamppost corona. Moreover, another reflection flavor relxillCp, in which the spectrum of the coronal emission is computed from the nthComp (Zdziarski et al., 1996) model of spherical geometry, also fits the data for Cyg X11-1- 1 (E6 and E7) and Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E4). Therefore, it is evident that the reflection spectroscopy also results in degeneracy among several models relying on contemporary coronal geometries.
@@
Overall, the wide-band spectro-temporal analysis (see §4.1 and §4.2) confirms the presence of distinct spectral states in BH-XRBs, but falls short of resolving the inherent degeneracy among the possible coronal geometries. This highlights the importance of incorporating X-ray polarimetric observations into the study of BH-XRBs. Subsequently, we explore how polarization measurements can offer deeper insights into this issue.
@@ 4.3 Polarization Measurements
We study the in-depth polarimetric properties of the eleven BH-XRBs observed with IXPE during multiple campaigns. In general, the polarization measurements are performed using two distinct methods. The first one relies on the model-independent approach based on the analysis of the polarization cube (Kislat et al., 2015) using the dedicated python package IXPEOBSSIMv31.0.3121212https://ixpeobssim.readthedocs.io/en/latest/ (Baldini et al., 2022). This provides the polarimetric properties by computing the polarization degree (PD) and polarization angle (PA) from the normalized Stokes parameters without any prior assumption on the emission mechanism. The alternative method depends on the simultaneous modeling of the I, Q, and U Stokes spectra in XPSEC. In this approach, the source intensity spectrum (I) is modeled using the appropriate physical components responsible for the emission mechanism and a polarization model accounting for the Q and U spectra (Strohmayer, 2017). Accordingly, we carry out both model-independent and model-dependent analyses to ensure the consistency of the polarimetric measurements.
@@
We perform model-independent polarization cube analysis using XPBIN tool inside the software routine with PCUBE algorithm. Following Kushwaha et al. 2023b; Majumder et al. 2024a, b, we consider all events from the three DUs of IXPE in the analysis. Subsequently, the normalized Q and U Stokes parameters, the polarization degree $({\rm PD}=\sqrt{(Q/I)^{2}+(U/I)^{2}})$ and polarization angle $({\rm PA}=\frac{1}{2}\tan^{-1}(U/Q))$ are computed in $2-8$ keV energy band for all observations of the respective sources. Following the guidelines of IAU [^13] https://aas.org/posts/news/2015/12/iau-calls-consistency-use-polarization-angle (Contopoulos & Jappel, 1974), we adopt the convention that PA increases counter-clockwise from North to East direction in the sky. We also calculate the minimum detectable polarization at $99\%$ confidence level ($\rm MDP_{99}\%$) and the significance (SIGNIF in units of $\sigma$) for each measurement. It may be noted that an observed $\rm PD>MDP_{99}\%$ indicates a polarization level that would arise from random statistical fluctuations with only $1\%$ chance probability (Kislat et al., 2015). All the calculated polarimetric parameters and the measurement statistics are tabulated in Table 5. In Fig. 4, we present $3\sigma$ confidence contours associated with the measurements of PD and PA for the respective sources in different spectral states.
@@
We find significant degree of polarized emission ($\rm PD>\rm MDP_{\rm 99}\%$) in all sources except LMC X11-1- 1, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 and MAXI J1744--294. More precisely, we find that PD varies within $1.22\pm 0.35\%-21.41\pm 0.41\%$ including all sources in different spectral states within $2-8$ keV energy band. We note that Cyg X33-3- 3 shows the maximum degree of polarization, whereas GX $339-4$ exhibits the lowest PD ($>\rm MDP_{99}\%$) among all sources under consideration. Further, a significant variation of PD is observed over different spectral states of the sources. For example, Cyg X11-1- 1 manifests a maximum PD of $\sim 4.8\%$ in the LHS, which decreases to $\sim 1.4\%$ in the HSS. Interestingly, a similar variation of PD is observed over several spectral states during the outburst phase of Swift J1727.816131727.816131727.8-16131727.8 - 1613 in which the polarization degree drops down to $\sim 0.4\%$ (HSS) from $\sim 4.7\%$ (HIMS) and again increases to $\sim 3.2\%$ in the LHS. Most astonishingly, 4U $1630-47$ exhibits exceedingly high PD in the HSS ($\sim 8.3\%$) and SPL ($\sim 6.8\%$), respectively. Moreover, we find an unprecedentedly high PD of $\sim 9\%$ in the LHS of the recent outburst of IGR J17091362417091362417091-362417091 - 3624, which is the maximum reported for any confirmed BH-XRB observed with IXPE to date (see also Ewing et al. 2025). In contrast, 4U $1957+115$ and LMC X33-3- 3 are found to show $\sim 1.9\%$ and $\sim 2.4-3\%$ polarization in the HSS, respectively. It may be noted that, for GX $339-4$, significant PD ($\sim 1.2\%$, E1) is obtained in $3-8$ keV energy range (see also Mastroserio et al. 2025), whereas a null-detection is seen considering the entire IXPE energy band of $2-8$ keV. Moreover, we find that the PA of the sources in different spectral states shows marginal variation (see Table 5).
@@
  • Null-detection of Polarization: It is worth mentioning that we re-confirm the null-detection of polarized emission ($\rm PD<\rm MDP_{\rm 99}\%$) for a few sources under consideration. In particular, LMC X11-1- 1 remains the first source that fails to manifest significant polarization with IXPE (see Table 5 and Podgorný et al. 2023). Further, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 exhibits a PD of $\sim 0.25\%$ in $2-8$ keV which remains significantly low from the corresponding $\rm MDP_{\rm 99}\sim 0.78\%$, confirming a null-detection. This finding contradicts the polarization detection with PD $\sim 1.3\%$ as reported by Mondal et al. (2024). This discrepancy possibly arises as the present analysis is carried out using the latest version of the software IXPEOBSSIMv31.0.3, fixing the bugs related to the visualization of polarization cubes in the previous releases [^14] https://ixpeobssim.readthedocs.io/en/latest/release_notes.html. We emphasize that our results are consistent with the recent polarization measurements of Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 Ling et al. (2024). Further, the null-detection of polarization is also observed in a few observations of Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E6, E7), GX $339-4$ (E2) and LMC X33-3- 3 (E3). In addition, for the first time, we report the null-detection of X-ray polarization with PD ($\sim 0.7\%$) $<\rm MDP_{\rm 99}(1.3\%)$ in the recently discovered BH-XRB candidate MAXI J174429417442941744-2941744 - 294 during HSS.

### v2

@@ 4.1.2 Power Density Spectra
We investigate the power density spectra (PDS) of all sources in different spectral states using NICER observations (see Table 1) in 0.5100.5-10 keV energy range. In addition, PDS properties in hard X-rays (>10>10 keV) are also examined using available NuSTAR and AstroSat observations of the sources. We generate 0.01 s time binned light curves and compute the PDS up to Nyquist frequency of $50$ Hz with $1024$ newbins per interval using powspec inside HEASOFT V6.34. The individual power spectra are averaged out to get the resultant PDS in units of $\rm rms^{2}/Hz$ which is geometrically rebinned in the frequency space with a factor of $1.03$ (see Belloni et al., 2002, 2005; Sreehari et al., 2019b; Majumder et al., 2022). In the upper panels of Fig. 2, we present the PDS of the respective sources in different spectral states. We model each PDS with a combination of constant and multiple Lorentzian components. Further, we compute the total percentage rms amplitude ($rms_{\rm tot}\%$) as a measure of variability in 0.1500.1-50 Hz and tabulate it in Table 3.
@@
Interestingly, we observe that the PDS in several spectral states exhibit distinct variable features. In particular, PDS of Cyg X1-1 in the LHS show significant variability with $rms_{\rm tot}\%\sim 31.1-52.4$ (0.5100.5-10 keV) and a clear power spectral break at $\sim 1.5$ Hz with NICER observations. Further, power spectral properties in hard X-rays (3803-80 keV) with quasi-simultaneous AstroSat observations exhibit similar break frequencies ($\sim 1.5$ Hz) with $rms_{\rm tot}\%\sim 37$ in LHS (see Table 1 and Fig. 2a). We also observe marginal (significant) PDS variability of $rms_{\rm tot}\%\sim 9.1$ ($33$) with low (high) energy NICER (AstroSat) observations of Cyg X1-1 in the HSS. Note that similar variability properties including power spectral breaks were also reported by Kushwaha et al. (2021) in LHS and HSS of Cyg X1-1.
@@
For Swift J1727.816131727.8-1613, the presence of type-C QPO feature at $\sim 1.4$ Hz is observed during HIMS (epoch 1, hereafter E1) in 0.51000.5-100 keV with NICER, NuSTAR and AstroSat quasi-simultaneous observations (see also Nandi et al. 2024; Liao et al. 2024). The evolution of the QPO frequency up to $\sim 6.7$ Hz is seen with NICER in the later epochs (E2-E5). However, a power distribution, similar to Cyg X1-1 (LHS), is observed in the LHS of Swift J1727.816131727.8-1613 without the presence of power spectral break, whereas marginal variability with $rms_{\rm tot}\%\sim 1.4$ is noticed during HSS (E6-E7).
@@
The remaining sources (4U 1630471630-47, LMC X1-1, LMC X3-3, 4U $1957+115$, Cyg X3-3 and Swift J151857.0572147151857.0-572147) in SIMS and HSS display weak variability signatures mostly represented by constant noise distribution and occasionally marginal power variation at lower frequencies. Conversely, during the recurrent outburst of GX 3394339-4, a sharp type-B QPO peak at $\sim 4.5$ Hz ($rms_{\rm QPO}\%\sim 10.2$) and harmonic feature at $\sim 9.5$ Hz along with weak broadband noise ($rms_{\rm tot}\%\sim 15.4$) are observed in SIMS (see Fig. 2 and Aneesha et al. 2024). Note that, the QPO features with enhanced $rms_{\rm QPO}\%$ from $3.76$ (0.5100.5-10 keV) to $12.6$ (3803-80 keV) are observed with NICER and AstroSat, respectively. For IGR J17091362417091-3624, a sharp type-C QPO at $\sim 0.2$ Hz ($rms_{\rm QPO}\%\sim 15$) along with a weak harmonic at $\sim 0.4$ Hz are observed during the recent outburst (see Fig. 2).
@@ 4.2.1 Spectral Modeling
Figure 3: Variation of the data to model ratio of the HSS/HIMS spectra of Cyg X1-1, GX 3394339-4 and Swift J1727.816131727.8-1613 fitted with model combination M1 and M3. The improvement in the respective fits by incorporating the reflection model component over M1 is depicted in panels (b), (d) and (f). See the text for details.
@@ 4.2.2 Alternative Spectral Models
[…]
@@ 4.2.3 Spectral Properties
Figure 4: Confidence contours ($3\sigma$) in the PD–PA plane obtained from the model-independent polarization measurements using 282-8 keV IXPE data, combining all DUs across various spectral states of the sources. The epoch of each observation is indicated in parentheses next to the corresponding spectral state. For Swift J151857.0572147151857.0-572147, LMC X1-1 and MAXI J17442941744-294, null-detection of polarization is observed, whereas all other sources exhibit significant polarization. Inset highlights a zoomed-in region of PD–PA space for better clarity.
@@
We find that the effects of Comptonization are noticeably high during the LHS of Cyg X-1 with covering fraction $cov_{\rm frac}\sim 0.61_{-0.04}^{+0.04}-0.77_{-0.02}^{+0.02}$, indicating Comptonization of up to $\sim 77\%$ of the soft photons in the LHS spectra. In addition, a high Comptonized flux contribution of $73-80\%$ is also observed in the spectra (see Table 3). Indeed, such a high $cov_{\rm frac}$ results in low photon index ($\Gamma_{\rm th}\lesssim 1.64_{-0.02}^{+0.02}$) as observed during the LHS observations of Cyg X1-1. In the HIMS and LHS of Swift J1727.8-1613, we find $cov_{\rm frac}\sim 0.35_{-0.01}^{+0.01}-0.82_{-0.02}^{+0.02}$ and $\Gamma_{\rm th}\sim 1.71_{-0.02}^{+0.03}-2.11_{-0.04}^{+0.05}$ with up to $\sim 64\%$ Comptonized emission. Similarly, LHS of IGR J17091-3624 which is dominated by the Comptonized emission ($\sim 87\%$ flux contribution) yields harder ($\Gamma_{\rm th}\sim 1.62_{-0.01}^{+0.02}$) spectral characteristics. On the other hand, the HSS/SIMS spectra of Cyg X-1 and Swift J1727.8-1613 results in low Comptonization with $cov_{\rm frac}$ in the range $0.005_{-0.002}^{+0.002}-0.38_{-0.03}^{+0.04}$ and thermal disc emission with $39-98\%$ flux contribution along with a steeper spectral index ($\Gamma_{\rm th}\sim 2.32_{-0.04}^{+0.06}-5.14_{-0.39}^{+0.36}$). Moreover, HSS with $76-98\%$ disc contribution are observed for LMC X-1, 4U 1630-47, LMC X-3 and 4U 1957++115 with negligible Comptonized emission ($cov_{\rm frac}<0.1$) as shown in Table 3. We find that GX 339-4 (Swift J151857.0-572147) exhibits intermediate spectral characteristics with $cov_{\rm frac}\sim 0.34$ ($0.21$) and $\Gamma_{\rm th}\sim 2$ ($2.5$) with $\sim 50\%$ ($78\%$) disc emission during SIMS (HSS).
@@
The reflection spectroscopy of Cyg X1-1, Swift J1727.816131727.8-1613 and GX 3394339-4 with model M3 during the HSS/SIMS results in disc ionization parameter ($\log\xi$) in the range of $3.31_{-0.05}^{+0.05}-4.29_{-0.21}^{+0.14}$ erg cm $\rm s^{-1}$, iron abundance ($A_{\rm fe}$) between $\sim 4.61_{-0.31}^{+0.36}$ and $6.35_{-1.46}^{+2.08}$ $A_{\odot}$, and reflection fraction ($R_{\rm f}$) ranging $\sim 0.59_{-0.11}^{+0.05}$ to $3.16_{-0.33}^{+0.35}$. Here, $A_{\odot}$ represents the solar iron abundance. The best-fitted parameters of the relxill component obtained from model M3 are tabulated in Table 4, whereas the parameters associated with other spectral components (diskbb and thcomp) of M3 are presented in Table 3 for the respective sources. Notably, a high ionization ($\log\xi\sim 4$) observed in HSS/SIMS generally suggests the reflection of significant coronal emission producing strong reflection signatures.
@@
In addition, we find that the model component relxilllp (Dauser et al., 2013) also delineates the reflection features assuming the primary source emissivity profile from a lamppost corona. Moreover, another reflection flavor relxillCp, in which the spectrum of the coronal emission is computed from the nthComp (Zdziarski et al., 1996) model of spherical geometry, also fits the data for Cyg X1-1 (E6 and E7) and Swift J1727.816131727.8-1613 (E4). Therefore, it is evident that the reflection spectroscopy also results in degeneracy among several models relying on contemporary coronal geometries.
@@
Overall, the wide-band spectro-temporal analysis (see §4.1 and §4.2) confirms the presence of distinct spectral states in BH-XRBs, but falls short of resolving the inherent degeneracy among the possible coronal geometries. This highlights the importance of incorporating X-ray polarimetric observations into the study of BH-XRBs. Subsequently, we explore how polarization measurements offer deeper insights to this issue.
@@ 4.3 Polarization Measurements
We study the in-depth polarimetric properties of the eleven BH-XRBs observed with IXPE during multiple campaigns. In general, the polarization measurements are performed using two distinct methods. The first one relies on the model-independent approach based on the analysis of the polarization cube (Kislat et al., 2015) using the dedicated python package IXPEOBSSIMv31.0.3777https://ixpeobssim.readthedocs.io/en/latest/ (Baldini et al., 2022). This provides the polarimetric properties by computing the polarization degree (PD) and polarization angle (PA) from the normalized Stokes parameters without any prior assumption on the emission mechanism. The alternative method depends on the simultaneous modeling of the I, Q, and U Stokes spectra in XPSEC. In this approach, the source intensity spectrum (I) is modeled using the appropriate physical components responsible for the emission mechanism and a polarization model accounting for the Q and U spectra (Strohmayer, 2017). Accordingly, we carry out both model-independent and model-dependent analyses to ensure the consistency of the polarimetric measurements.
@@
We perform model-independent polarization cube analysis using XPBIN tool inside the software routine with PCUBE algorithm. Following Kushwaha et al. 2023; Majumder et al. 2024a, b, we consider all events from the three DUs of IXPE in the analysis. Subsequently, the normalized Q and U Stokes parameters, the polarization degree $({\rm PD}=\sqrt{(Q/I)^{2}+(U/I)^{2}})$ and polarization angle $({\rm PA}=\frac{1}{2}\tan^{-1}(U/Q))$ are computed in 282-8 keV energy band for all observations of the respective sources. Following the guidelines of IAU [^8] https://aas.org/posts/news/2015/12/iau-calls-consistency-use-polarization-angle (Contopoulos & Jappel, 1974), we adopt the convention that PA increases counter-clockwise from North to East direction in the sky. We also calculate the minimum detectable polarization at $99\%$ confidence level ($\rm MDP_{99}\%$) and the significance (SIGNIF in units of $\sigma$) for each measurement. It may be noted that an observed $\rm PD>MDP_{99}\%$ indicates a polarization level that would arise from random statistical fluctuations with only $1\%$ chance probability (Kislat et al., 2015). All the calculated polarimetric parameters and the measurement statistics are tabulated in Table 5. In Fig. 4, we present $3\sigma$ confidence contours associated with the measurements of PD and PA for the respective sources in different spectral states.
@@
We find significant degree of polarized emission ($\rm PD>\rm MDP_{\rm 99}\%$) in all sources except LMC X1-1, Swift J151857.0572147151857.0-572147 and MAXI J1744-294. More precisely, we find that PD varies within $1.22\pm 0.35\%-21.41\pm 0.41\%$ including all sources in different spectral states within 282-8 keV energy band. We note that Cyg X3-3 shows the maximum degree of polarization, whereas GX 339-4 exhibits the lowest PD ($\rm MDP_{99}\%$) among all sources under consideration. Further, a significant variation of PD is observed over different spectral states of the sources. For example, Cyg X1-1 manifests a maximum PD of $\sim 4.8\%$ in the LHS, which decreases to $\sim 1.4\%$ in the HSS. Interestingly, a similar variation of PD is observed over several spectral states during the outburst phase of Swift J1727.816131727.8-1613 in which the polarization degree drops down to $\sim 0.4\%$ (HSS) from $\sim 4.7\%$ (HIMS) and again increases to $\sim 3.2\%$ in the LHS. Most astonishingly, 4U 1630471630-47 exhibits exceedingly high PD in the HSS ($\sim 8.3\%$) and SPL ($\sim 6.8\%$), respectively. Moreover, we find an unprecedentedly high PD of $\sim 9\%$ in the LHS of the recent outburst of IGR J17091362417091-3624, which is the maximum reported for any confirmed BH-XRB observed with IXPE to date (see also Ewing et al. 2025). In contrast, 4U 1957++115 and LMC X-3 are found to show $\sim 1.9\%$ and $\sim 2.4-3\%$ polarization in the HSS, respectively. It may be noted that, for GX 339-4, significant PD ($\sim 1.2\%$, E1) is obtained in 3-8 keV energy range (see also Mastroserio et al., 2025), whereas null-detection is seen considering the entire IXPE energy band of 2-8 keV. Moreover, we find that PA of the sources in different spectral states shows marginal variation (see Table 5).
@@
Null-detection of Polarization: It is worth mentioning that we re-confirm the null-detection of polarized emission ($\rm PD<\rm MDP_{\rm 99}\%$) for a few sources under consideration. In particular, LMC X1-1 remains the first source that fails to manifest significant polarization with IXPE (see Table 5 and Podgorný et al. 2023). Further, Swift J151857.0572147151857.0-572147 exhibits a PD of $\sim 0.25\%$ in 282-8 keV which remains significantly low from the corresponding $\rm MDP_{\rm 99}\sim 0.78\%$, confirming a null-detection. This finding contradicts the polarization detection with PD $\sim 1.3\%$ as reported by Mondal et al. (2024). This discrepancy possibly arises as the present analysis is carried out using the latest version of the software IXPEOBSSIMv31.0.3, fixing the bugs related to the visualization of polarization cubes in the previous releases [^9] https://ixpeobssim.readthedocs.io/en/latest/release_notes.html. We emphasize that our results are consistent with the recent polarization measurements of Swift J151857.0572147151857.0-572147 Ling et al. (2024). Further, the null-detection of polarization is also observed in a few observations of Swift J1727.816131727.8-1613 (E6, E7), GX 3394339-4 (E2) and LMC X3-3 (E3). In addition, for the first time, we report the null-detection of X-ray polarization with PD ($\sim 0.7\%$) $\rm MDP_{\rm 99}(1.3\%)$ in the recently discovered BH-XRB candidate MAXI J17442941744-294 during HSS.